OpenStudy (ray1998):

Help? Will give fan and medal. What is the vertex form of the equation? y = -x^2 + 12x - 4

OpenStudy (solomonzelman):

do you know what a "perfect trinomial" is?

OpenStudy (ray1998):

I'm sorry! I got busy for a little bit. And although it sounds very familiar, I can't remember. @SolomonZelman

OpenStudy (solomonzelman):

There is a formula; \(\color{#000000 }{ \displaystyle (x+a)^2=x^2+2ax+a^2 }\) And the right side is called a "perfect square trinomial" because the right side is a perfect square, except that it is a trinomial - not a number. When you multiply \(\color{#000000 }{ \displaystyle (x+a)}\) times another \(\color{#000000 }{ \displaystyle (x+a) }\), you get; \(\color{#000000 }{ \displaystyle (x+a)\cdot (x+a)= }\) \(\color{#000000 }{ \displaystyle x(x+a)+ a(x+a)= }\) \(\color{#000000 }{ \displaystyle x^2+ax+ ax+a^2= }\) \(\color{#000000 }{ \displaystyle x^2+2ax+a^2. }\) (that proves the formula, and it is a perfect square. Just as when you multiply 4•4 and get 16, 16 is a perfect square because it's square root is an integer 4, SO IS \(\color{#000000 }{ \displaystyle x^2+2ax+a^2 }\) a perfect square, because when you multiply (x+a) times (x+a) you get \(\color{#000000 }{ \displaystyle x^2+2ax+a^2 }\)) So, we can infer that for a trinomial to be a "perfect square" we have to put it in a form \(\color{#000000 }{ \displaystyle x^2+2ax+a^2 }\). You might wonder why you need to know about it, but the vertex form of the parabola looks like this: \(\color{#000000 }{ \displaystyle y=a(x-h)^2+k }\) And notice we have \(\color{#000000 }{ \displaystyle (x-h)^2 }\) sitting in there. Aha! This is why we need a perfect square trinomial, because \(\color{#000000 }{ \displaystyle x^2+2ax+a^2 }\) is (same as) \(\color{#000000 }{ \displaystyle (x+a)^2 }\); that way we can put our equation into the "vertex form". \(\color{red}{_\text{___________________________________________________________________}}\) More about vertex form eq. of parabola: Form; \(\color{#000000 }{ \displaystyle y=a(x-h)^2+k }\) Vertex; \(\color{#000000 }{ \displaystyle (h,k) }\) Opens down: \(\color{#000000 }{ \displaystyle a<0 }\) (negative a) vertex is the maximum point (maximum of y=k, that occurs at x=h) Opens up: \(\color{#000000 }{ \displaystyle a>0 }\) (positive a) vertex is the minimum point (minimum of y=k, that occurs at x=h) \(\color{red}{^\text{___________________________________________________________________}}\)

OpenStudy (ray1998):

It makes sense :) thank you very much!

OpenStudy (ray1998):

This was really helpful.

OpenStudy (solomonzelman):

site glitches...

OpenStudy (solomonzelman):

I am putting an example up

OpenStudy (ray1998):

Site glitches?

OpenStudy (solomonzelman):

I will try to give an example to clarify some things... Suppose I want to write \(\color{#000000 }{ \displaystyle y=x^2-4x }\) in a vertex form. Then, I will need to put my equation in a form of \(\color{#000000 }{ \displaystyle x^2+2ax+a^2 }\). I already have the \(\color{#ff0000 }{ \displaystyle x^2+2ax }\) (and that is \(\color{#000000 }{ \displaystyle x^2-4x }\)), so what I need to do, is to add "\(a^2\)" component, and that I will be able to do only if I figure out the \(a\). I can find \(a\) via comparison. \(\color{#ff0000 }{ \displaystyle x^2+2ax }\) in the form, CORRESPONDS to \(\color{#ff0000 }{ \displaystyle x^2-4x }\) in our equation, so, we can tell that \(2a=-4\) --> \(a=-2\). Therefore, "\(a^2\)" in our case is going to be \(\color{#000000 }{ \displaystyle a^2\Longrightarrow (-2)^2=4 }\) So we would need \(\color{#000000 }{ \displaystyle y=x^2-4x\color{blue}{+4} }\) but, we can NOT just stick a \(4\) in there (right?) So, we will do a trick! \(\color{#000000 }{ \displaystyle y=x^2-4x\color{blue}{+4}\color{red}{-4} }\) (by adding 4 and subtracting 4, I am not changing the value) Now, we can see the perfect square part underlined, \(\color{#000000 }{ \displaystyle y=\underline{x^2-4x\color{blue}{+4}}\color{red}{-4} }\) and this therefore becomes, \(\color{#000000 }{ \displaystyle y=\underline{(x-2)^2}\color{red}{-4} }\) (I have -2 in the parenthesis, because I already found a to be -2) So \(\color{#000000 }{ \displaystyle y=x^2-4x }\) is written as \(\color{#000000 }{ \displaystyle y=(x-2)^2-4 }\) in a vertex form.

OpenStudy (solomonzelman):

(if you want another example or have questions.... u know....)